# Beamforming Design for Intelligent Reflecting Surface Aided Near-Field THz Communications

### v1

@@ III Proposed IRS Beamforming
Accordingly, the normailized piece-wise far-field cascaded BS-IRS-user channel between the BS and the user, through the reflection of the $(s_{y},s_{z})$-th IRS element in the $(k_{y},k_{z})$-th sub-surface, can be denoted by $\tilde{\mathbf{b}}_{m}$, whose entries are given by
$\left[\tilde{\mathbf{b}}_{m}\right]^{k_{y},k_{z}}_{s_{y},s_{z}}=$ $e^{j2\pi\frac{f_{m}}{c}\left(r^{br}_{k_{y},k_{z}}-r^{ru}_{k_{y},k_{z}}\right)}$
×$\times e^{-j2\pi\frac{f_{m}}{c}\left(\varphi^{br,k_{y},k_{z}}_{s_{y},s_{z}}-\varphi^{ru,k_{y},k_{z}}_{s_{y},s_{z}}\right)},\vspace{-0.3cm}$ (13)
where
$\varphi^{br,k_{y},k_{z}}_{s_{y},s_{z}}=$ $\Delta_{s_{y}}^{S}d\sin\left(\psi^{ele}_{br,k_{y},k_{z}}\right)\sin\left(\psi^{azi}_{br,k_{y},k_{z}}\right)$
$+\Delta^{S}_{s_{z}}dr^{br}_{k_{y},k_{z}}\cos\left(\psi^{ele}_{br,k_{y},k_{z}}\right)$ (14)
and
$\varphi^{ru,k_{y},k_{z}}_{s_{y},s_{z}}=$ $\Delta_{s_{y}}^{S}d\sin\left(\psi^{ele}_{ru,k_{y},k_{z}}\right)\sin\left(\psi^{azi}_{ru,k_{y},k_{z}}\right)$
$+\Delta^{S}_{s_{z}}dr^{ru}_{k_{y},k_{z}}\cos\left(\psi^{ele}_{ru,k_{y},k_{z}}\right).$ (15)
@@
The piece-wise far-field channel model makes it straightforward to decouple the phase in (III) into two components: the inter-subsurface phase difference $\Delta_{k_{y},k_{z}}^{r}=r^{br}_{k_{y},k_{z}}-r^{ru}_{k_{y},k_{z}}$ among different sub-surfaces, and the intra-subsurface phase difference $\Delta\varphi^{k_{y},k_{z}}_{s_{y},s_{z}}=\varphi^{br,k_{y},k_{z}}_{s_{y},s_{z}}-\varphi^{ru,k_{y},k_{z}}_{s_{y},s_{z}}$ within each sub-surface. Considering the fact that each sub-surface contains much fewer elements, this reduces the effect of beam split. Hence, we can reasonably deduce that the intra-surface phase difference $\Delta\varphi^{k_{y},k_{z}}_{s_{y},s_{z}}$ has little effect on the near-field beam split. Thus, we design the TD values to compensate for the inter-surface phase difference $\Delta_{k_{y},k_{z}}^{r}$ in order to alleviate the near-field beam split effect. Accordingly, the required phase shift at the $(k_{y},k_{z})$-th sub-surface to combat beam split effect at the $m$-th subcarrier is given as
@@
$\tau_{k_{y},k_{z}}=-\frac{\Delta_{k_{y},k_{z}}^{r}}{c},\forall k_{y},k_{z}.$ (17)
@@
However, the delay value can be large in this case. For example, with $\Delta_{k_{y},k_{z}}^{r}=-3$ meters (m), $\tau_{k_{y},k_{z}}$ can be $10000$ ps. To avoid excess TD module’s limitation, we propose a DLDD architecture as shown in Fig. 3. The first layer TD network utilizes $\left(K_{y}-1\right)$ TD modules in a step-wise manner, being able to generate $K_{y}$ cumulative TD values, each of which is connected to a group of $\left(K_{z}-1\right)$ TD modules in the second layer TD network. Thus, the DLDD architecture contains a total number of $K_{t}=\left(K_{y}-1\right)+K_{y}\times\left(K_{z}-1\right)=\left(K-1\right)$ TD modules. Specifically, the $k_{y}^{t}$-th delay value, with $k_{y}^{t}=\{1,\dots,K_{y}-1\}$, in the first layer TD network, is designed to eliminate the phase difference between the $(k_{y}^{t},1)$-th and the $(k_{y}^{t}+1,1)$-th sub-surfaces, which is given as
$\Delta\tau_{k_{y}^{t}}^{f}=\tau_{k_{y}^{t}+1,1}-\tau_{k_{y}^{t},1}=-\frac{\Delta_{k_{y}^{t}+1,1}^{r}-\Delta_{k_{y}^{t},1}^{r}}{c},\forall k_{y}^{t}.\vspace{-0.15cm}$ (18)
@@
Similarly, the $k_{z}^{t}$-th delay value, with $k_{z}^{t}=\{1,\dots,K_{z}-1\}$, of the $k_{y}$-th group in the second layer TD network is designed to eliminate the phase difference between the $(k_{y},k_{z}^{t}+1)$-th and the $(k_{y},k_{z}^{t})$-th sub-surfaces, which is given as
$\Delta\tau_{k_{y},k_{z}^{t}}^{s}=\tau_{k_{y},k_{z}^{t}+1}-\tau_{k_{y},k_{z}^{t}}=-\frac{\Delta_{k_{y},k_{z}^{t}+1}^{r}-\Delta_{k_{y},k_{z}^{t}}^{r}}{c},\forall k_{y},k_{z}^{t}.$ (19)
@@
In this way, the delay value applied on the $(k_{y},k_{z})$-th sub-surface is realized by a series of TD modules, i.e., $\tau_{k_{y},k_{z}}=\sum_{k=1}^{k_{y}-1}\Delta\tau_{k}^{f}+\sum_{i=1}^{k_{z}-1}\Delta\tau_{k_{y},i}^{s}$, rather than by a dedicated TD module as in [6]. Through such cumulative TDs, the proposed DLDD IRS beamforming architecture alleviates the maximum delay limitation of TD modules. The TD vector for the first layer TD network is expressed as $\mathbf{t}^{\Delta}_{f}=[\Delta\tau_{1}^{f},\dots,\Delta\tau_{K_{y}-1}^{f}]$, and that for the second layer TD network is expressed as $\mathbf{t}^{\Delta}_{s}=[\Delta\tau_{1,1}^{s},\Delta\tau_{1,2}^{s}\dots,\Delta\tau_{K_{y},K_{z}-1}^{s}]$. As demonstrated in (18) and (19), the TD values can be either positive or negative. However, TD modules are unable to generate negative delays. Fortunately, we have the following proposition.
Proposition 1
The sign of $\tau_{k_{y},k_{z}},\forall k_{y},k_{z}$ is consistent, and the same holds true for $\Delta\tau_{k_{y}^{t}}^{f},\forall k_{y}^{t}$ and $\Delta\tau_{k_{y},k_{z}^{t}}^{s},\forall k_{y},k_{z}^{t}$.
@@
It is proved that the IRS needs to be deployed near the BS or the user for optimal performance [12], which yields $r_{br}\gg r_{ru}$ or $r_{br}\ll r_{ru}$. It is intuitive that, in either case, the sign of $\tau_{k_{y},k_{z}},\forall k_{y},k_{z}$ is consistent. Taking $r_{br}\ll r_{ru}$ as an example, as $k_{y}$ increases, it is obvious that the increase in $r^{br}_{k_{y},k_{z}}$ is more significant than that in $r^{ru}_{k_{y},k_{z}}$, which results in negative sign of $\Delta\tau_{k_{y}^{t}}^{f},\forall k_{y}^{t}$, and vice verse. The consistency of the sign of $\Delta\tau_{k_{y},k_{z}^{t}}^{s},\forall k_{y},k_{z}^{t}$ can be similarly proved. This completes the proof. ∎
Building upon proposition 1, we propose to insert a $2$-output switch before each group of TD modules. The switch closes its $1^{st}$ output when the elements of $\mathbf{t}^{\Delta}_{f}$ or $\mathbf{t}^{\Delta}_{s}$ are positive, or rather directs the signal through the $2^{nd}$ output to the TD network.
@@
On the other hand, the IRS phase shift is designed to align with the intra-surface phase discrepancy $\Delta\varphi^{k_{y},k_{z}}_{s_{y},s_{z}}$. Specifically, for the $(s_{y},s_{z})$-th IRS element in the $(k_{y},k_{z})$-th sub-surface, we have
$\boldsymbol{\theta}^{k_{y},k_{z}}_{s_{y},s_{z}}=\operatorname{mod}\left(2\pi\frac{f_{c}}{c}\Delta\varphi^{k_{y},k_{z}}_{s_{y},s_{z}},2\pi\right),\forall k_{y},k_{z},s_{y},s_{z},$ (20)
where $\operatorname{mod}$ is the modulo operation. To this end, the design of the proposed DLDD IRS beamforming scheme is achieved by (18), (19) and (20).
IV Numerical Results

### v2

@@ III-C Proposed DLDD IRS Beamforming
Accordingly, the normalized piecewise far-field cascaded BS-IRS-user channel between the BS and the user, through the reflection of the $(s_{y},s_{z})$-th IRS element in the $(k_{y},k_{z})$-th sub-surface, can be denoted by $\tilde{\mathbf{b}}_{m}$, whose entries are given by
$\vspace{-0.3cm}\left[\tilde{\mathbf{b}}_{m}\right]^{k_{y},k_{z}}_{s_{y},s_{z}}=$ $e^{-j2\pi\frac{f_{m}}{c}\left(r_{k_{y},k_{z}}+l_{k_{y},k_{z}}\right)}$
×$\times e^{j2\pi\frac{f_{m}}{c}\left(\varphi^{k_{y},k_{z}}_{s_{y},s_{z}}+\varepsilon^{k_{y},k_{z}}_{s_{y},s_{z}}\right)},$ (14)
@@
where $\varphi^{k_{y},k_{z}}_{s_{y},s_{z}}\!=\!\Delta_{s_{y}}^{S}d\sin\!\left(\!\psi_{k_{y},k_{z}}\!\right)\!\sin\!\left(\!\phi_{k_{y},k_{z}}\!\right)\!+\!\Delta^{S}_{s_{z}}d\cos\!\left(\!\psi_{k_{y},k_{z}}\!\right)\!$ and $\varepsilon^{k_{y},k_{z}}_{s_{y},s_{z}}\!=\!\Delta_{s_{y}}^{S}d\!\sin\!\left(\!\upsilon_{k_{y},k_{z}}\!\right)\!\sin\!\left(\!\omega_{k_{y},k_{z}}\!\right)\!+\!\Delta^{S}_{s_{z}}d\cos\!\left(\!\upsilon_{k_{y},k_{z}}\!\right)\!$. The piecewise far-field channel model makes it straightforward to decouple the phase in (III-C) into two components: the inter-subsurface phase difference $\delta_{k_{y},k_{z}}^{r}=r_{k_{y},k_{z}}+l_{k_{y},k_{z}}$ among different sub-surfaces, and the intra-subsurface phase difference $\delta^{k_{y},k_{z}}_{s_{y},s_{z}}=\varphi^{k_{y},k_{z}}_{s_{y},s_{z}}+\varepsilon^{k_{y},k_{z}}_{s_{y},s_{z}}$ within each sub-surface. Considering the fact that each sub-surface contains much fewer elements, this reduces the effect of beam split. Hence, we can reasonably deduce that the intra-surface phase difference $\delta^{k_{y},k_{z}}_{s_{y},s_{z}}$ has little effect on the near-field beam split. Thus, we design the TD values to compensate for the inter-surface phase difference $\delta_{k_{y},k_{z}}^{r}$ in order to alleviate the near-field beam split effect. Accordingly, the required phase shift at the $(k_{y},k_{z})$-th sub-surface to combat beam split effect at the $m$-th subcarrier is given by
@@
$\tau_{k_{y},k_{z}}=-\frac{\delta_{k_{y},k_{z}}^{r}}{c},k_{y}\in\mathcal{K}_{y},k_{z}\in\mathcal{K}_{z}.$ (16)
@@
However, the delay value can be large in this case. For example, with $\delta_{k_{y},k_{z}}^{r}=3$ meters (m), $\tau_{k_{y},k_{z}}$ can be $1\times 10^{4}$ ps. To avoid excess TD module’s limitation, we propose a DLDD architecture as shown in Fig. 3, which utilizes a two-layer network configuration. The first layer TD network consists of $\left(K_{y}-1\right)$ TD modules arranged in a step-wise manner, generating $K_{y}$ cumulative delays. Each of these cumulative delays is then fed into an array of $\left(K_{z}-1\right)$ TD modules. As a result, the second layer consists of a total of $K_{y}\left(K_{z}-1\right)$ TD modules. Therefore, the proposed DLDD architecture requires a total of $K_{\text{T}}=\left(K_{y}-1\right)+K_{y}\times\left(K_{z}-1\right)=\left(K-1\right)$ TD modules.
@@
The first-layer TD network is specifically designed to eliminate the phase difference between adjacent sub-surfaces along the $y$-axis. Specifically, the $k_{y}$-th delay value in this layer, $k_{y}\in\mathcal{K}_{y}^{\prime}\triangleq\{1,\dots,K_{y}-1\}$, is designed to eliminate the phase difference between the $(k_{y},1)$-th and the $(k_{y}+1,1)$-th sub-surfaces, which is given by
$\Delta\tau_{k_{y}}^{\text{f}}\!=\!\tau_{k_{y}+1,1}\!-\!\tau_{k_{y},1}\!=\!-\frac{\delta_{k_{y}+1,1}^{r}\!-\!\delta_{k_{y},1}^{r}}{c},k_{y}\in\mathcal{K}_{y}^{\prime}.$ (17)
Similarly, the second-layer TD network aims at eliminating the phase difference between adjacent sub-surfaces along the $z$-axis. Specifically, the $k_{z}$-th delay value, $k_{z}\in\mathcal{K}_{z}^{\prime}\triangleq\{1,\dots,K_{z}-1\}$, of the $k_{y}$-th group is designed to eliminate the phase difference between the $(k_{y},k_{z}+1)$-th and the $(k_{y},k_{z})$-th sub-surfaces, which is given by
$\Delta\tau_{k_{y},k_{z}}^{s}$ $=\tau_{k_{y},k_{z}+1}-\tau_{k_{y},k_{z}}$
$=-\frac{\delta_{k_{y},k_{z}+1}^{r}-\delta_{k_{y},k_{z}}^{r}}{c},k_{y}\in\mathcal{K}_{y},k_{z}\in\mathcal{K}_{z}^{\prime}.$ (18)
Thus, the total delay applied on the $(k_{y},k_{z})$-th sub-surface is achieved through a cumulative sum of TD values from both layers, expressed as $t_{k_{y},k_{z}}=\sum_{k=1}^{k_{y}-1}\Delta\tau_{k}^{\text{f}}+\sum_{i=1}^{k_{z}-1}\Delta\tau_{k_{y},i}^{\text{s}}$.
@@
Such cumulative strategy allows a more efficient utilization of TD modules, as the architecture does not rely on dedicated TD modules for each delay value, significantly alleviating the maximum delay constraints faced by conventional systems [6]. To represent the design mathematically, the parameters of the TD modules in the first layer network is expressed as $\boldsymbol{\tau}_{\text{f}}=[\Delta\tau_{1}^{\text{f}},\dots,\Delta\tau_{K_{y}-1}^{\text{f}}]$, and that in the second layer TD network is expressed as $\boldsymbol{\tau}_{s}=[\Delta\tau_{1,1}^{\text{s}},\Delta\tau_{1,2}^{\text{s}}\dots,\Delta\tau_{K_{y},K_{z}-1}^{\text{s}}]$. As demonstrated in (17) and (III-C), the TD values can be either positive or negative. However, TD modules are unable to generate negative delays. Fortunately, we have the following Lemma.
Lemma 1. Define $\epsilon\left(k_{y}\right)\triangleq-\frac{\left(y_{\text{b}}-\Delta_{k_{y}}^{K_{y}}Sd\right)}{r_{k_{y},1}}-\frac{\left(y_{\text{u}}-\Delta_{k_{y}}^{K_{y}}Sd\right)}{l_{k_{y},1}}$. TD values $\Delta\tau_{k_{y}}^{\text{f}}$ are strictly positive, i.e., $\Delta\tau_{k_{y}}^{\text{f}}>0,\forall k_{y}\in\mathcal{K}_{y}^{\prime}$, if $\epsilon\left(K_{y}\right)<0$. Conversely, TD values are strictly negative, i.e., $\Delta\tau_{k_{y}}^{\text{f}}<0,\forall k_{y}\in\mathcal{K}_{y}^{\prime}$, if $\epsilon\left(1\right)>0$.
@@
The first-order derivative of $\delta_{k_{y},1}^{r}$ w.r.t. $k_{y}$ is given by $\frac{d(\delta_{k_{y},1}^{r})}{d\left(k_{y}\right)}=\epsilon\left(k_{y}\right)$. In addition, since the second-order derivative of $\delta_{k_{y},1}^{r}$ w.r.t. $k_{y}$ satisfies $\frac{d^{2}(\delta_{k_{y},1}^{r})}{d\left({k_{y}}\right)^{2}}=\frac{x_{\text{b}}^{2}+\left(z_{\text{b}}-\Delta_{1}^{K_{z}}Sd\right)^{2}}{r_{k_{y},1}}+\frac{x_{\text{u}}^{2}+\left(z_{\text{u}}-\Delta_{1}^{K_{z}}Sd\right)^{2}}{l_{k_{y},1}}\!>\!0,\forall k_{y}$, it indicates the monotonic increase in $\epsilon_{k_{y}}$. If $\epsilon\left(K_{y}\right)<0$, then $\epsilon\left(k_{y}\right)<0,\forall k_{y}\leq K_{y}$, which readily indicates that $\Delta\tau_{k_{y}}^{\text{f}}\!\!>\!\!0,\forall k_{y}\in\mathcal{K}_{y}^{\prime}$. Conversely, if $\epsilon\left(1\right)>0$, then $\epsilon\left(k_{y}\right)>0,\forall k_{y}\geq 1$, implying $\Delta\tau_{k_{y}}^{\text{f}}\!\!<\!\!0,\forall k_{y}\in\mathcal{K}_{y}^{\prime}$. The proof is thus complete. ∎
Remark 1
The property of $\Delta\tau_{k_{y},k_{z}}^{s}$ can be similarly proved. Thus, we focus on the discussion on $\Delta\tau_{k_{y}}^{\text{f}}$, which can be directly applied to $\Delta\tau_{k_{y},k_{z}}^{s}$.
Lemma 1 establishes a connection between the property of the TD sequence $\Delta\tau_{k_{y}}^{\text{f}}$ and the locations of the transceivers. Since the size of the IRS is generally far less than the distances $r_{k_{y},1}$ and $l_{k_{y},1}$, we reasonably assume that $\epsilon_{k_{y}}>0$ for all $k_{y}\in\mathcal{K}_{y}$ if $\epsilon_{i}>0$ for any $i\in\mathcal{K}_{y}$, and vice versa. In other words, it is reasonable to assume that $\epsilon_{1}>0$ and $\epsilon_{K_{y}}<0$ never occur simultaneously. Under such assumption, the signs of TD values $\Delta\tau_{k_{y}}^{\text{f}},\forall k_{y}\in\mathcal{K}_{y}^{\prime}$ are consistent. To this end, we propose to insert a $2$-output switch before each group of TD modules. The switch closes its $1^{st}$ output when the elements of $\boldsymbol{\tau}_{\text{f}}$ or $\boldsymbol{\tau}_{s}$ are positive, or rather directs the signal through the $2^{nd}$ output to the TD network. Specifically, it is modeled as a binary selecting vector $\mathbf{s}_{\text{\text{f}}}$, given by
$\mathbf{s}_{\text{\text{f}}}=\frac{1}{2|\Delta\tau_{1}^{\text{f}}|}\left[\begin{array}[]{l}|\Delta\tau_{1}^{\text{f}}|+\Delta\tau_{1}^{\text{f}}\\ |\Delta\tau_{1}^{\text{f}}|-\Delta\tau_{1}^{\text{f}}\end{array}\right].$ (19)
Denote $\mathbf{t}^{+}_{\text{\text{f}}}=\left[0,\Delta\tau_{1}^{\text{f}},\sum_{k_{y}=1}^{2}\Delta\tau_{k_{y}}^{\text{f}}\cdots,\sum_{k_{y}=1}^{K_{y}-1}\Delta\tau_{k_{y}}^{\text{f}}\right]$ and $\mathbf{t}^{-}_{\text{f}}=\left[\sum_{k_{y}=1}^{K_{y}-1}\Delta\tau_{k_{y}}^{\text{f}},\sum_{k_{y}=2}^{K_{y}-1}\Delta\tau_{k_{y}}^{\text{f}},\cdots,\Delta\tau_{K_{y}-1}^{f},0\right]$. The resulting delay of the first layer TD network is then given by
$\mathbf{t}_{\text{f}}=\left[\mathbf{t}^{+}_{\text{f}},\mathbf{t}^{-}_{\text{f}}\right]\mathbf{s}_{\text{f}}.\vspace{-0.2cm}$ (20)
@@
On the other hand, the IRS phase shift is designed to align with the intra-surface phase discrepancy $\delta^{k_{y},k_{z}}_{s_{y},s_{z}}$. Specifically, for the $(s_{y},s_{z})$-th IRS element in the $(k_{y},k_{z})$-th sub-surface, we have
$\theta^{k_{y},k_{z}}_{s_{y},s_{z}}=\operatorname{mod}\left(-2\pi\frac{f_{c}}{c}\delta^{k_{y},k_{z}}_{s_{y},s_{z}},2\pi\right),\forall k_{y},k_{z},s_{y},s_{z},$ (21)
where $\operatorname{mod}$ is the modulo operation. Denote $t_{\text{f},k_{y}}$ and $t_{\text{s},k_{y},k_{z}}$ as the corresponding entries of $\mathbf{t}_{\text{f}}$ and $\mathbf{t}_{\text{s},k_{y}}$, where $\mathbf{t}_{\text{s},k_{y}}$ is the resulting delay of the $k_{y}$-th group in second layer TD network, derived similarly as above. Thus, the total delay applied on the $(k_{y},k_{z})$-th sub-surface is re-expressed as $t_{k_{y},k_{z}}=t_{\text{f},k_{y}}+t_{\text{s},k_{y},k_{z}}$. With the above design, the array gain is expressed as (22), shown at the top of the page.
$\eta\left(f_{m}\right)=\left|\sum_{k_{y}=1}^{K_{y}}\sum_{k_{z}=1}^{K_{z}}e^{-j2\pi f_{m}t_{k_{y},k_{z}}}\sum_{s_{y}=1}^{S}\sum_{s_{z}=1}^{S}e^{j2\pi\frac{f_{m}}{c}\left(r^{k_{y},k_{z}}_{s_{y},s_{z}}+l^{k_{y},k_{z}}_{s_{y},s_{z}}\right)}e^{j\theta^{k_{y},k_{z}}_{s_{y},s_{z}}}\right|.\vspace{-0.4cm}$ (22)
IV Extension to Joint Beamforming for Solving Double-Stage Beam Split Effect
In this section, we extend our discussion from single-antenna to multi-antenna at the BS, where beam split occurs at both the IRS and the BS. In fact, the near-field double-stage beam split is coupled, making an optimal beamforming solution unattainable. Building on a similar approach as in the previous section, we propose a sub-optimal joint beamforming scheme to address this issue.
Again, we start by analyzing the distance of between the BS and the IRS. Following segmentation strategy in the previous section, we divide the entire BS antenna array into $K_{t}$ sub-arrays, each of which contains $N_{k}$ adjacent antennas, i.e., $N_{t}=K_{t}\times N_{k}$. Then, we can approximate the distance between the $n_{k}$-th BS antenna in the $k_{t}$-th sub-array and the $(s_{y},s_{z})$-th element of $(k_{y},k_{z})$-th sub-surface as
$\vspace{-0.3cm}r^{n_{k},k_{t}}_{(s_{y},s_{z}),(k_{y},k_{z})}\approx$ $r^{k_{t}}_{k_{y},k_{z}}-\Delta_{s_{y}}^{S_{y}}d\sin\left(\psi^{k_{t}}_{k_{y},k_{z}}\right)\sin\left(\phi^{k_{t}}_{k_{y},k_{z}}\right)$
$-\Delta_{s_{z}}^{S_{z}}$ $\cos\left(\psi^{k_{t}}_{k_{y},k_{z}}\right)-\Delta_{n_{k}}^{N_{k}}d\cos\left(\psi^{k_{t}}_{k_{y},k_{z}}\right),\vspace{-0.3cm}$ (23)
where $r^{k_{t}}_{k_{y},k_{z}}$ is the distance between the center of $k_{t}$-th BS sub-array and the center of $(k_{y},k_{z})$-th IRS sub-surface, while $\phi^{k_{t}}_{k_{y},k_{z}}$ and $\psi^{k_{t}}_{k_{y},k_{z}}$ denote the corresponding azimuth and elevation angles, satisfying $r^{k_{t}}_{k_{y},k_{z}}\!=\!\sqrt{x_{\text{b}}^{2}\!+\!\left(y_{\text{b}}\!-\!\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}\!\!+\!\left(z_{\text{b}}\!-\!\Delta_{k_{t}}^{K_{t}}N_{k}d\!-\!\Delta_{k_{z}}^{K_{z}}Sd\right)^{2}}$, $\sin\!\left(\!\phi^{k_{t}}_{k_{y},k_{z}}\!\right)\!\!=\!\!\sin\!\left(\!\phi_{k_{y},k_{z}}\!\right)$, $\sin\!\left(\!\psi^{k_{t}}_{k_{y},k_{z}}\!\right)\!\!=\!\!\frac{\left(x_{\text{b}}^{2}+\left(y_{\text{b}}-\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}\right)^{\frac{1}{2}}}{r^{k_{t}}_{k_{y},k_{z}}}$, and $\cos\!\left(\!\psi^{k_{t}}_{k_{y},k_{z}}\!\right)\!\!=\!\!\frac{z_{\text{b}}-\Delta_{k_{t}}^{K_{t}}N_{k}d-\Delta_{k_{z}}^{K_{z}}Sd}{r^{k_{t}}_{k_{y},k_{z}}}$. Then, the channel between the BS and the IRS can be approximated based on (IV), denoted by $\tilde{\mathbf{G}}_{m}$.
Motivated by the decomposability of the far-field double-stage beam split proved in [7], where the beam split at each stage is addressed by the IRS beamforming and BS beamforming separately, we propose a sub-optimal joint beamforming scheme. Recall that the proposed DLDD IRS beamforming strategy in section III-C aims at solving the phase discrepancy of between the signal reflected by different IRS elements, i.e., the entries in one of the columns in $\tilde{\mathbf{G}}_{m}$. Thus, we design the BS beamforming to mitigate the phase discrepancy along one of the rows in $\tilde{\mathbf{G}}_{m}$. The design of BS beamforming is given by
$\vspace{-0.1cm}\mathbf{w}_{m}=\mathbf{P}\mathbf{d}_{m}=\operatorname{blkdiag}\left(\left[\mathbf{p}_{1},\cdots,\mathbf{p}_{K_{t}}\right]\right)e^{j2\pi\frac{f_{m}}{c}\mathbf{t}},$ (24)
where $\mathbf{p}_{k_{t}}$ is the phase shift at the $k_{t}$-th sub-array, designed as
$\mathbf{p}_{k_{t}}\!\!=\!\!\left[\!e^{\!-j2\pi\frac{f_{c}}{c}\Delta_{1}^{N_{k}}d\cos\left(\!\psi^{k_{t}}_{k_{y},k_{z}}\!\right)}\!,\!\cdots\!,\!e^{\!-j2\pi\frac{f_{c}}{c}\Delta_{N_{k}}^{N_{k}}d\cos\left(\!\psi^{k_{t}}_{k_{y},k_{z}}\!\right)}\!\right]^{T}\!\!,$ (25)
and $\mathbf{t}$ is the resulting TD over the BS antenna array, designed as
$\vspace{-0.1cm}\mathbf{t}=\left[0,\tau_{1},\cdots,\sum_{i=1}^{K_{t}-1}\tau_{i}\right],$ (26)
where $\tau_{i}=r^{i}_{1,1}-r^{i+1}_{1,1}$. To this end, we have provided the design of the BS beamforming. Together with the IRS beamforming, such joint beamforming scheme solves the double-stage beam split effect sub-optimally.
V Numerical Results